1 Given a market demand curve is Qd = 200 – 2P and the equation for a market supply curve is Qs = 50 + 3P, calculate the market equilibrium price and quantity. (5)


1
Expert's answer
2021-09-16T09:57:36-0400

"1. Equilibrium price"


"200-2P = 50+3P"


"200-50 = 3P + 2P"


"150 = 5P"


"P = \frac{150}{5} = 30"


Equilibrium price = 30


"2. Equilibrium quantity"


"Qd = 200-2P"


"= 200-2(30)"


"= 200-60 = 140"


Equilibrium quantity = 140